25. For each of the linear operators below, determine whether it is normal, self adjoint, or neither (a) T : R2 → R2 defined by T(x, y) = (2x − 2y, −2x + 5y). (b) T : C 2 → C 2 defined by T(x, y) = (2x + iy, x + 2y). 


1
Expert's answer
2022-04-15T03:32:11-0400

"a:\\\\T=\\left[ \\begin{matrix}\t2&\t\t-2\\\\\t-2&\t\t5\\\\\\end{matrix} \\right] \\\\T^*=\\left[ \\begin{matrix}\t2&\t\t-2\\\\\t-2&\t\t5\\\\\\end{matrix} \\right] =T\\\\TT^*=T^2=TT^*\\\\Normal, self-adjoint\\\\b:\\\\T=\\left[ \\begin{matrix}\t2&\t\ti\\\\\t1&\t\t2\\\\\\end{matrix} \\right] \\\\T^*=\\left[ \\begin{matrix}\t2&\t\t1\\\\\t-i&\t\t2\\\\\\end{matrix} \\right] \\ne T\\\\TT^*=\\left[ \\begin{matrix}\t2&\t\ti\\\\\t1&\t\t2\\\\\\end{matrix} \\right] \\left[ \\begin{matrix}\t2&\t\t1\\\\\t-i&\t\t2\\\\\\end{matrix} \\right] =\\left[ \\begin{matrix}\t5&\t\t2+2i\\\\\t2-2i&\t\t5\\\\\\end{matrix} \\right] \\\\TT^*=\\left[ \\begin{matrix}\t2&\t\t1\\\\\t-i&\t\t2\\\\\\end{matrix} \\right] \\left[ \\begin{matrix}\t2&\t\ti\\\\\t1&\t\t2\\\\\\end{matrix} \\right] =\\left[ \\begin{matrix}\t5&\t\t2+2i\\\\\t2-2i&\t\t5\\\\\\end{matrix} \\right] =TT^*\\\\Normal, not\\,\\,self-adjoint"